# HyperSHAP: Shapley Values and Interactions for Hyperparameter Importance

### v1

@@ A.3 Example: Non-Monotone Sensitivity Game
and set the baseline HPC to $\bm{\lambda}^{0}:=(0,0)$. The Sensitivity game values are then computed by observing that $\textsc{Val}_{u}(\bm{\lambda}^{\ast})$ with $\bm{\lambda}^{\ast}\sim p^{\ast}(\bm{\lambda}^{\ast})$ is described as a Bernoulli variable Ber($q$). For $S=\{1,2\}$, the probability of $\textsc{Val}_{u}$ being 1 is $q=1/4$, since both hyperparameters have to be set to zero. In contrast, for $|S|=1$, we have $q=1/2$, since the remaining variable is already set at zero due to the baseline HPC. We thus obtain again the variances with $q(1-q)$ as
@@ Appendix B Experimental Setup
Our implementation builds upon the shapiq package (Muschalik et al., 2024a), which is publicly available on GitHub [^1] https://github.com/mmschlk/shapiq, for computing Shapley values and interactions. Furthermore, for the experiments, we use YAHPO-Gym (Pfisterer et al., 2022), a surrogate-based benchmark for multi-fidelity hyperparameter optimization. YAHPO-Gym provides several benchmark suites, i.a., lcbench (Zimmer et al., 2021), which we focused on in the main paper. However, in the subsequent sections, we also present results from the rbv2_ranger benchmark set, a random forest benchmark, from YAHPO-Gym demonstrating the more general applicability of HyperSHAP. Furthermore, we run evaluations on the benchmark PD1 and JAHS-Bench-201 to showcase HyperSHAP’s wide applicability. In our repository, we provide pre-computed games to foster reproducibility of our results and allow for faster post-processing of the game values, e.g., for plotting different representations of the played games.
@@ B.1 Considered Benchmarks
lcbench (Pfisterer et al., 2022; Zimmer et al., 2021).
lcbench is a benchmark considering joint optimization of the neural architecture and hyperparameters that has been proposed by (Zimmer et al., 2021) together with the automated deep learning system Auto-PyTorch. The benchmark consists of 35 datasets with 2000 configurations each for which the learning curves have been recorded, allowing for benchmarking multi-fidelity HPO. However, in YAHPO-Gym only 34 of the 35 original datasets are contained which is why our evaluation is also restricted to those 34 datasets.
@@
rbv2_ranger (Pfisterer et al., 2022).
@@
JAHS-Bench-201 (Bansal et al., 2022).
@@ B.2 Approximation of the argmax
As per Definition 4.3 to Definition 4.6, for every coalition $S$, we need to determine the $\arg\max$. However, the true $\arg\max$ is difficult to determine, so we approximate it throughout our experiments. For the sake of implementation simplicity and unbiased sampling, we use random search with a large evaluation budget of 10 0001000010\,00010 000 candidate evaluations. As the configurations are independently sampled, for evaluating a configuration, we simply blind an initially sampled batch of 10,000 hyperparameter configurations for the hyperparameters not contained in the coalition $S$ by setting their values to the default value. This procedure is fast to compute and reduces the noise potentially occurring through randomly sampling entirely new configurations for every coalition evaluation. After blinding, the surrogate model provided by YAHPO-Gym is then queried for the set of hyperparameter configurations, and the maximum observed performance is returned.
@@
Figure 7: Hyperparameter importance with HyperSHAP, approximating the $\arg\max$ in Definition 4.3 of the value function via hyperparameter optimization with increasing budgets for dataset ID 7593 of lcbench. For tuning, we consider the following hyperparameters of lcnet: learning rate (L-R), batch size (B-S), weight decay (W-D), num layers (N-L), momentum (M), max units (M-U), and max dropout (M-D).
@@ B.4 Hardware Usage and Compute Resources
Initial computations for lcbench and rbv2_ranger have been conducted on consumer hardware, i.e., Dell XPS 15 (Intel i7 13700H, 16GB RAM) and a MacBook Pro (M3 Max - 16C/40G, 128GB RAM). Overall computations took around 10 CPUd, highlighting HyperSHAP being lightweight when combined with surrogates. In the course of the reviewing process, we re-computed the games for Ablation and Data-Specific Multi-Data Tunability of lcbench and rbv2_ranger and added PD1 and JAHS-Bench-201. These computations have been conducted on a high-performance computer with nodes equipped with 2$\times$ AMD Milan 7763 ($2\times 64$ cores) and 256GiB RAM of which 1 core and 8GB RAM have been allocated to the computations for a single game. While the latter experiments amounted to 10.71 CPU days, in sum, the computations for this paper accumulate roughly 21 CPU days. The average runtimes per benchmark and game are as follows (Table 2):
@@
PD1 4 4 64.9$\pm$16.0 862.4$\pm$13.7 -
JAHS 10 3 123.7$\pm$4.4 30,406.7$\pm$4750.9 (8h26m) -
LCBench 7 34 4.8$\pm$0.4 357.3$\pm$3.1 10,713.4 (2h58m)
rbv2_ranger 8 119 26.4$\pm$6.8 6,717$\pm$767.3 -
@@ Appendix C Guidance on Interpreting Interaction Visualizations
To visualize and interpret lower-, and higher-order interactions such as SI or MI, we employ the SI graph visualization and the UpSet plot from shapiq (Muschalik et al., 2024a). The SI graph visualization is an extension of the network plot for Shapley interactions (Muschalik et al., 2024b) and can be used to visualize higher order interactions. The UpSet plot (Lex et al., 2014) is a well-established method for visualizing set-based scores, which can also be used for representing higher-order interactions. Figure 8 shows an exemplary SI graph and UpSet plot.
@@
An UpSet plot for SIs or MIs shows a selection of high-impact interactions and their scores. The plot is divided into two parts. The upper part shows the interaction values as bars and the lower part shows the considered interactions as a matrix. The first two bars in Figure 8 show the main effects of the O-M and L-I hyperparameters. The third bar shows the negative interaction of both of these features (denoted as the connection between the interactions). A red color denotes a positive score, and a blue color denotes a negative score. The bars and interactions are plotted in descending order according to the absolute value of an interaction (i.e., higher-impact interactions first).
@@
An SI graph plot in Figure 8 can be interpreted as follows. Each individual player (e.g. hyperparameter) is represented as a node with connecting hyperedges representing the strength and direction of interactions. Akin to the well-established force plots (Lundberg & Lee, 2017), positive interactions are colored in red and negative interactions in blue, respectively. The strength of an interaction is represented by the size and opacity of the hyperedge. To reduce visual clutter, small interactions below a predefined absolute threshold may be omitted from the graph. Notably, first-order interactions (i.e., individual player contributions, or main effects) are represented by the size of the nodes.
Figure 8: An UpSet plot (left) and a SI graph plot (right) for the Tunability game from Section 5.1.
@@ D.1 Measuring the Magnitude of Interactions
In this section, we provide further details for measuring the presence of interactions discussed in Section 5.2. The MIs describe the pure additive effect of a coalition to the payout of the game. They thus serve as an important tool to analyze the interactions present in a game $\nu$. For instance, low-complexity games, where MIs are non-zero only up to coalitions of size $k$, are typically referred as $k$-additive games (Grabisch, 2016). In this case, SIs with explanation order $k$ perfectly recover all game values (Bordt & von Luxburg, 2023). In this case, the SIs correspond to the MIs. We thus analyze the absolute values of MIs for varying size of coalitions, i.e., displaying the strata $q(k):=\{|m(S)|:S\subseteq\mathcal{N},|S|=k\}$ for varying interaction order $k=1,\dots,n$. Analyzing $q(k)$ indicates, if the game $\nu$ has lower- order higher-order interactions present by investigating the magnitudes and distributions in the strata $q(k)$.
@@ D.2 Analyzing Lower-Order Representations of Games
In this section, we provide additional details for the lower-order representations and $R^{2}$ scores discussed in Section 5.2. The SV that capture the fair contribution in a game $\nu$ of an individual to the joint payout $\nu(\mathcal{N})$. However, the SV $\phi^{\text{SV}}(i)$ is also the solution to a constrained weighted least squares problem (Charnes et al., 1988; Fumagalli et al., 2024b)
@@
In other words, the SV is the best additive approximation of the game $\nu$ in terms of this weighted loss constrained on the efficiency axiom. Based on this result, the FSII (Tsai et al., 2023) was introduced as
@@
where the infinite weights capture the constraints $\nu(\emptyset)=\Phi_{k}(\emptyset)$ and $\nu(\mathcal{N})=\sum_{S\subseteq\mathcal{N}}\Phi_{k}(S)$. Note that Tsai et al. (2023) introduce FSII with a scaled variant of $\mu$ that does not affect the solution. The FSII can thus be viewed as the best possible approximation of the game $\nu$ using additive components up to order $k$ constrained on the efficiency axiom. It is therefore natural to introduce the Shapley-weighted faithfulness as
@@
In our experiments, we rely on FSII, since this interaction index optimizes the faithfulness measure $\mathcal{F}$ by definition. However, $k$-Shapley Value ($k$-SII) satisfies a similar faithfulness property (Fumagalli et al., 2024b). Since the FSII is equal to the MIs for $k=n$, we have that $\mathcal{F}(\nu,\Phi_{n})=0$ due to the additive recovery property of the MIs. Hence, $R^{2}(n)=R^{2}(\nu,\Phi_{n})=0$ in this case. Clearly, the $R^{2}(k)$ scores are monotonic increasing in $k$ by definition of FSII. An $R^{2}(k)\approx 1$ indicates an almost perfect recovery of all game values. In our experiments, we have shown that higher-order interactions are present, but lower-order representations (low $k$) are mostly sufficient to achieve very high $R^{2}$ scores. This indicates that higher-order interactions are present but do not dominate the interaction landscape in our applications. For instance, a single isolated higher-order interaction would yield much lower $R^{2}$ scores (Muschalik et al., 2024a).
@@ D.3 Additional Experimental Details
In Section 5.2, we investigate how faithful HyperSHAP explanations capture the interaction structures of the HPO problem. For this we compute Tunability explanations for all four benchmarks, lcbench, rbv2_ranger PD1, and JAHS-Bench-201. Further, we compute Multi-Data Tunability explanations for lcbench and rbv2_ranger over all instances in the benchmarks. We then compute the MIs for all of these explanations. We compute HyperSHAP FSII explanations up to the highest order. Then we compute the Shapley-weighted $R^{2}$ loss between the explanations and the original game as a measure of faithfulness. Figure 9 summarizes the results. The high $R^{2}$ score (almost $1.0$) for both the Tunability and the Multi-Data Tunability games suggests that most of the explanatory power is captured by interactions up to the third order, confirming prior research that suggests hyperparameter interactions are typically of lower order (Pushak & Hoos, 2020).
Figure 9: Detailed Reprint of Figure 3 (right). Curves for Multi-Data Tunability contain only one game each. The Tunability games for lcbench and rbv2_ranger are averaged over $20$ randomly selected datasets. The Tunability curves for PD1 and JAHS-Bench-201 are averaged over all datasets contained in the benchmarks ($4$ and $3$, respectively). The shaded bands correspond to the standard error of the mean (SEM).
@@ Appendix E Additional Empirical Results
This section contains additional experimental results, including more detailed plots and visualizations for the experiments conducted in Section 5.
@@ E.1 Additional Information for the Comparison of Ablation and Tunability
In Section 5.1, we compare the Ablation and the Tunability settings and see that we can derive different interpretations from both explanations into the Hyperparameter optimization. Interpreting the Ablation explanation suggests that only the lr_initial (L-I) hyperparameter is important for achieving high performance. However, the Tunability explanation reveals that actually both, the opt_momentum (O-M) and initial_learning_rate L-I, hyperparameters are useful for tuning. The optimizer needs to decide which hyperparameter to focus on. Figure 10 contains shows the same result as in Figure 2 with more detail.
@@
Figure 10: UpSet (left) and SI graph (right) plots for the Ablation (top) and Tunability (bottom) settings described in Section 5.1. The SI graph plots show all interactions and the UpSet plot the ten most impactful interactions.
@@ E.2 Additional Results for Comparison with fANOVA
This section contains additional results for the evaluation of hyperparameter optimization runs restricted to the top-2 important hyperparameters according to fANOVA (Hutter et al., 2014), Sensitivity, and Tunability of HyperSHAP. Figure 11 shows that selecting and tuning hyperparameters with HyperSHAP leads to better anytime performance than with fANOVA or Sensitivity. The suggested top-2 hyperparameters for every method are listed in Table 3. We can observe that overall, although not always perfect, HyperSHAP suggests a top-2 that yields higher anytime performance, meaning that the hyperparameter optimizer achieves a higher accuracy quicker. However, hyperparameters are suggested with respect to their overall hyperparameter importance, which does not necessarily guarantee better anytime performance as these hyperparameters can be more difficult to tune than others with lower impact. Still, in this case, the lower impact hyperparameters could result in better anytime performance for smaller budgets. We consider an in-depth study of which hyperparameters to suggest for which subsequent HPO task to be an interesting avenue of future work.

### v2

@@ A.4 Example: Non-Monotone Sensitivity Game
and set the baseline configuration to $\bm{\lambda}^{0}:=(0,0)$. The Sensitivity game values are then computed by observing that $\textsc{Val}_{u}(\bm{\lambda}^{\ast})$ with $\bm{\lambda}^{\ast}\sim p^{\ast}(\bm{\lambda}^{\ast})$ is described as a Bernoulli variable Ber($q$). For $S=\{1,2\}$, the probability of $\textsc{Val}_{u}$ being 1 is $q=1/4$, since both hyperparameters have to be set to zero. In contrast, for $|S|=1$, we have $q=1/2$, since the remaining variable is already set at zero due to the baseline configuration. We thus obtain again the variances with q(1q)q(1-q) as
@@ Appendix B Experimental Setup
Our implementation builds upon the shapiq package (version 1.1.1) (Muschalik.2024), which is publicly available on GitHub [^1] https://github.com/mmschlk/shapiq and distributed via pypi, for computing Shapley values and interactions. Furthermore, for the experiments, we use YAHPO-Gym (DBLP:conf/automl/PfistererSMBB22), a surrogate-based benchmark for multi-fidelity hyperparameter optimization. YAHPO-Gym provides several benchmark suites, i.a., lcbench (zimmer-tpami21a), which we focused on in the main paper. However, in the subsequent sections, we also present results from the rbv2_ranger benchmark set, a random forest benchmark, from YAHPO-Gym, demonstrating the more general applicability of HyperSHAP. Furthermore, we run evaluations on the benchmark PD1 and JAHS-Bench-201 to showcase HyperSHAP’s wide applicability. In our repository, we provide pre-computed games to foster reproducibility of our results and allow for faster post-processing of the game values, e.g., for plotting different representations of the played games.
For better readability in terms of the font size, hyperparameter names are abbreviated in the interaction graphs. An overview of abbreviations and the names of their corresponding hyperparameters can be found in the subsequent section.
@@ B.1 Considered Benchmarks
lcbench (DBLP:conf/automl/PfistererSMBB22; zimmer-tpami21a).
lcbench is a benchmark considering joint optimization of the neural architecture and hyperparameters that has been proposed by (zimmer-tpami21a) together with the automated deep learning system Auto-PyTorch. The benchmark consists of 35 datasets with 2000 configurations each for which the learning curves have been recorded, allowing for benchmarking multi-fidelity HPO. However, in YAHPO-Gym only 34 of the 35 original datasets are contained which is why our evaluation is also restricted to those 34 datasets.
@@
rbv2_ranger (DBLP:conf/automl/PfistererSMBB22).
@@
JAHS-Bench-201 (bansal-neurips22a).
@@ B.2 Approximation of the argmax
As per definition 3 to definition 4, for every coalition $S$, we need to determine the $\arg\max$. However, the true $\arg\max$ is difficult to determine, so we approximate it throughout our experiments. For the sake of implementation simplicity and unbiased sampling, we use random search with a large evaluation budget of 10 00010\,000 candidate evaluations. As the configurations are independently sampled, for evaluating a configuration, we simply blind an initially sampled batch of 10,000 hyperparameter configurations for the hyperparameters not contained in the coalition $S$ by setting their values to the default value. This procedure is fast to compute and reduces the noise potentially occurring through randomly sampling entirely new configurations for every coalition evaluation. After blinding, the surrogate model provided by YAHPO-Gym is then queried for the set of hyperparameter configurations, and the maximum observed performance is returned.
@@
Figure 7: Hyperparameter importance with HyperSHAP, approximating the $\arg\max$ in Definition 3 of the value function via hyperparameter optimization with increasing budgets for dataset ID 7593 of lcbench. For tuning, we consider the following hyperparameters of lcnet: learning rate (L-R), batch size (B-S), weight decay (W-D), num layers (N-L), momentum (M), max units (M-U), and max dropout (M-D).
@@ B.4 Hardware Usage and Compute Resources
[…]
@@
[…]
After vectorization, implementing parallelization, and leveraging approximation methods for 15 hyperparmeters (HPs) and more, HyperSHAP’s runtimes for Tunability become much faster (https://github.com/automl/hypershap):
Table 3: Mean runtimes in seconds across 10 runs of HyperSHAP’s package implementation for Tunability. Values that have not been computed are indicated by ”-”.
#Cores / #HPs 2 3 4 5 6 7 8 9 10 11 12 13 14 15
1 0.2 0.3 0.5 1.0 2.0 3.9 7.8 15.8 31.7 62.7 125.0 251.0 501.4 -
8 - - - 0.3 0.5 0.9 1.8 3.5 6.9 13.9 27.4 55.2 111.4 97.2
@@ Appendix C Guidance on Interpreting Interaction Visualizations
To visualize and interpret lower-, and higher-order interactions such as SI or MI, we employ the SI graph visualization and the UpSet plot from shapiq (Muschalik.2024). The SI graph visualization is an extension of the network plot for Shapley interactions (DBLP:conf/aaai/MuschalikFHH24) and can be used to visualize higher order interactions. The UpSet plot (Lex2014UpSet) is a well-established method for visualizing set-based scores, which can also be used for representing higher-order interactions. Fig. 8 shows an exemplary SI graph and UpSet plot.
@@
An UpSet plot for SIs or MIs shows a selection of high-impact interactions and their scores. The plot is divided into two parts. The upper part shows the interaction values as bars and the lower part shows the considered interactions as a matrix. The first two bars in Fig. 8 show the main effects of the O-M and L-I hyperparameters. The third bar shows the negative interaction of both of these features (denoted as the connection between the interactions). A red color denotes a positive score, and a blue color denotes a negative score. The bars and interactions are plotted in descending order according to the absolute value of an interaction (i.e., higher-impact interactions first).
@@
An SI graph plot in Fig. 8 can be interpreted as follows. Each individual player (e.g. hyperparameter) is represented as a node with connecting hyperedges representing the strength and direction of interactions. Akin to the well-established force plots (Lundberg.2017), positive interactions are colored in red and negative interactions in blue, respectively. The strength of an interaction is represented by the size and opacity of the hyperedge. To reduce visual clutter, small interactions below a predefined absolute threshold may be omitted from the graph. Notably, first-order interactions (i.e., individual player contributions, or main effects) are represented by the size of the nodes.
Figure 8: An UpSet plot (left) and a SI graph plot (right) for the Tunability game from Sec. 6.1.
@@ D.1 Measuring the Magnitude of Interactions
In this section, we provide further details for measuring the presence of interactions discussed in Sec. 6.2. The MIs describe the pure additive effect of a coalition to the payout of the game. They thus serve as an important tool to analyze the interactions present in a game $\nu$. For instance, low-complexity games, where MIs are non-zero only up to coalitions of size $k$, are typically referred as $k$-additive games (Grabisch.2016). In this case, SIs with explanation order $k$ perfectly recover all game values (Bord.2023). In this case, the SIs correspond to the MIs. We thus analyze the absolute values of MIs for varying size of coalitions, i.e., displaying the strata $q(k):=\{|m(S)|:S\subseteq\mathcal{N},|S|=k\}$ for varying interaction order $k=1,\dots,n$. Analyzing $q(k)$ indicates, if the game $\nu$ has lower- order higher-order interactions present by investigating the magnitudes and distributions in the strata $q(k)$.
@@ D.2 Analyzing Lower-Order Representations of Games
In this section, we provide additional details for the lower-order representations and $R^{2}$ scores discussed in Sec. 6.2. The SV that capture the fair contribution in a game $\nu$ of an individual to the joint payout $\nu(\mathcal{N})$. However, the SV $\phi^{\text{SV}}(i)$ is also the solution to a constrained weighted least squares problem (Charnes.1988; DBLP:conf/icml/FumagalliMKHH24)
@@
In other words, the SV is the best additive approximation of the game $\nu$ in terms of this weighted loss constrained on the efficiency axiom. Based on this result, the FSII (Tsai.2022) was introduced as
@@
where the infinite weights capture the constraints $\nu(\emptyset)=\Phi_{k}(\emptyset)$ and $\nu(\mathcal{N})=\sum_{S\subseteq\mathcal{N}}\Phi_{k}(S)$. Note that (Tsai.2022) introduces FSII with a scaled variant of $\mu$ that does not affect the solution. The FSII can thus be viewed as the best possible approximation of the game $\nu$ using additive components up to order $k$ constrained on the efficiency axiom. It is therefore natural to introduce the Shapley-weighted faithfulness as
@@
In our experiments, we rely on FSII, since this interaction index optimizes the faithfulness measure $\mathcal{F}$ by definition. However, $k$-Shapley Value ($k$-SII) satisfies a similar faithfulness property (DBLP:conf/icml/FumagalliMKHH24). Since the FSII is equal to the MIs for $k=n$, we have that $\mathcal{F}(\nu,\Phi_{n})=0$ due to the additive recovery property of the MIs. Hence, $R^{2}(n)=R^{2}(\nu,\Phi_{n})=0$ in this case. Clearly, the $R^{2}(k)$ scores are monotonic increasing in $k$ by definition of FSII. An $R^{2}(k)\approx 1$ indicates an almost perfect recovery of all game values. In our experiments, we have shown that higher-order interactions are present, but lower-order representations (low $k$) are mostly sufficient to achieve very high $R^{2}$ scores. This indicates that higher-order interactions are present but do not dominate the interaction landscape in our applications. For instance, a single isolated higher-order interaction would yield much lower $R^{2}$ scores (Muschalik.2024).
@@ D.3 Additional Experimental Details
In Sec. 6.2, we investigate how faithful HyperSHAP explanations capture the interaction structures of the HPO problem. For this we compute Tunability explanations for all four benchmarks, lcbench, rbv2_ranger PD1, and JAHS-Bench-201. Further, we compute Multi-Data Tunability explanations for lcbench and rbv2_ranger over all instances in the benchmarks. We then compute the MIs for all of these explanations. We compute HyperSHAP FSII explanations up to the highest order. Then we compute the Shapley-weighted $R^{2}$ loss between the explanations and the original game as a measure of faithfulness. Fig. 9 summarizes the results. The high $R^{2}$ score (almost $1.0$) for both the Tunability and the Multi-Data Tunability games suggests that most of the explanatory power is captured by interactions up to the third order, confirming prior research that suggests hyperparameter interactions are typically of lower order (DBLP:conf/gecco/PushakH20a).
Figure 9: Detailed Reprint of Fig. 2 (right). Curves for Multi-Data Tunability contain only one game each. The Tunability games for lcbench and rbv2_ranger are averaged over 2020 randomly selected datasets. The Tunability curves for PD1 and JAHS-Bench-201 are averaged over all datasets contained in the benchmarks (44 and 33, respectively). The shaded bands correspond to the standard error of the mean (SEM).
@@ Appendix E Additional Empirical Results
This section contains additional experimental results, including more detailed plots and visualizations for the experiments conducted in Sec. 6.
@@ E.1 Additional Information for the Comparison of Ablation and Tunability
In Sec. 6.1, we compare the Ablation and the Tunability settings and see that we can derive different interpretations from both explanations into the Hyperparameter optimization. Interpreting the Ablation explanation suggests that only the lr_initial (L-I) hyperparameter is important for achieving high performance. However, the Tunability explanation reveals that actually both, the opt_momentum (O-M) and initial_learning_rate L-I, hyperparameters are useful for tuning. The optimizer needs to decide which hyperparameter to focus on. Fig. 10 contains shows the same result as in Fig. 3 with more detail.
@@
Figure 10: UpSet (left) and SI graph (right) plots for the Ablation (top) and Tunability (bottom) settings described in Sec. 6.1. The SI graph plots show all interactions and the UpSet plot the ten most impactful interactions.
@@ E.2 Additional Results for Comparison with fANOVA
This section contains additional results for the evaluation of hyperparameter optimization runs restricted to the top-2 important hyperparameters according to fANOVA (fANOVA), Sensitivity, and Tunability of HyperSHAP. Fig. 11 shows that selecting and tuning hyperparameters with HyperSHAP leads to better anytime performance than with fANOVA or Sensitivity. The suggested top-2 hyperparameters for every method are listed in Tab. 4. We can observe that overall, although not always perfect, HyperSHAP suggests a top-2 that yields higher anytime performance, meaning that the hyperparameter optimizer achieves a higher accuracy quicker. However, hyperparameters are suggested with respect to their overall hyperparameter importance, which does not necessarily guarantee better anytime performance as these hyperparameters can be more difficult to tune than others with lower impact. Still, in this case, the lower impact hyperparameters could result in better anytime performance for smaller budgets. We consider an in-depth study of which hyperparameters to suggest for which subsequent HPO task to be an interesting avenue of future work.